# The Cayley-Bacharach property and the Levinson-Ullery conjecture

### v1

@@ Abstract.
In this paper, we study the geometric configurations of a finite set of points having the Cayley-Bacharach property in the $n$-dimensional projective space ${\mathbb{P}}^{n}$. Our main contribution is the establishment of the Levinson-Ullery conjecture for the previously unsolved case where $d=4$ and all $r\geq 1$.
@@ 1. Introduction
Inspired by this result, in [16], Levinson and Ullery introduced the notion of a plane configuration to describe the geometry of $\mathbb{X}$} and proposed a conjecture below. Here a plane configuration $\mathcal{P}$ is a union of distinct positive-dimensional linear spaces $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ in ${\mathbb{P}}^{n}$, its dimension is $\dim(\mathcal{P})=\sum_{i=1}^{k}\dim(\mathcal{P}_{i})$ and its length is $l(\mathcal{P})=k$.
@@ 2. Preliminary
    The image of the vanishing ideal $I_{\mathbb{Y}}\subseteq P$ of $\mathbb{Y}$} in $R=P/I_{\mathbb{X}}$ is denoted by $I_{{\mathbb{Y}}/{\mathbb{X}}}=I_{\mathbb{Y}}/I_{\mathbb{X}}$. Its initial degree is $\alpha_{{\mathbb{Y}}/{\mathbb{X}}}:=\min\{k\in{\mathbb{N}}\mid(I_{{\mathbb{Y}}/{\mathbb{X}}})_{k}\neq 0\}$. Then $\alpha_{{\mathbb{Y}}/{\mathbb{X}}}\leq r_{\mathbb{X}}$ and the Hilbert function of $\mathbb{Y}$} satisfies
@@
Now we recall from [16, Definition 2.1] the following notion.
@@
Note that a 0-dimensional plane configuration is empty, a 1-dimensional plane configuration is a single line, and a 2-dimensional plane configuration is either two lines or one 2-plane. Here a $d$-plane refers to a $d$-dimensional linear spaces. Also, the linear span of a plane configuration $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ is denoted by $\mathop{\rm span}\nolimits(P)$ or $\mathop{\rm span}\nolimits(\mathcal{P}_{i}\mid i=1,\dots,k)$.
@@ 3. Proof of main theorem
    AA and $L$ are split. In this case, ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{B}$ are $\mathop{\rm CBP}\nolimits(r)$ by Proposition 2.8(a). Due to Corollary 2.4 and (a), the lower bounds for the cardinalities of ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{B}$ are given by
@@
    This implies $d(r+1)+4\leq|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|=|{\mathbb{X}}|\leq(d+1)r+1$, which is impossible.
@@
    AA and $L$ intersect at a point, say $p$. By Lemma 2.8(c), at least one of ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{A}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r).$ If ${\mathbb{X}}_{A}$ has $\mathop{\rm CBP}\nolimits(r)$, then claim (a) yields that $|{\mathbb{X}}_{A}|\geq 4r+2$. When ${\mathbb{X}}_{A}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r),$ then $|{\mathbb{X}}_{A}\cup\{p\}|\geq 4r+2$ by (a). In these both cases, we always have $|{\mathbb{X}}_{A}|\geq 4r+1$. On the other hand, the set ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$, and Corollary 2.4 shows that $|{\mathbb{X}}_{B}|\geq r+1$. Hence, we obtain $|{\mathbb{X}}|=|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|\geq(4r+1)+(r+1)\geq 5r+2$, a contradiction.
@@ 1. Case 1:
        If $|{\mathbb{X}}_{B}|=2r\leq 3(r-1)+1$, then ${\mathbb{X}}_{B}$ lies on a plane configuration of dimension 2 and $|{\mathbb{X}}_{A}|=3r+1$. When ${\mathbb{X}}_{B}$ lies on a line, as above, the set $\mathbb{X}$} lies on a plane configuration of dimension 4 and we are done. If ${\mathbb{X}}_{B}$ lies on a union of two split lines, then each of these lines contains at least $r+1$ points of ${\mathbb{X}}_{B}$ by Proposition 2.8(a) and Corollary 2.4, and so $|{\mathbb{X}}_{B}|\geq 2r+2$, which is impossible. Here we remark that a skew plane configuration of two lines agrees with a split plane configuration of two lines. Now we consider the last subcase that the set ${\mathbb{X}}_{B}$ lies on a $2$-plane $H^{\prime}$. If $H^{\prime}\cap H\neq\emptyset$ or $H^{\prime}\cap L\neq\emptyset$ then the pair $(H,H^{\prime})$ defines a 4-plane containing at least $4r$ points of $\mathbb{X}$} or the pair $(L,H^{\prime})$ and a point in ${\mathbb{X}}_{A}\cap H$ define a 4-plane containing at least $3r+2$ points of $\mathbb{X}$}, which contradicts to the choice of the 4-plane $A$. Hence we must have $H^{\prime}\cap H=H^{\prime}\cap L=\emptyset$. In particular, $\mathbb{X}$} is contained in a skew plane configuration $\mathcal{P}=H\cup L\cup H^{\prime}$ of length $\ell(\mathcal{P})=3<r$. Since $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$, an application of Proposition 2.8(b) yields that the line $L$ contains at least $r+2$ points of $\mathbb{X}$}. It follows that
@@ 1. Case 2:
    Since ${\mathbb{X}}_{B}$ is not contained in a $4$-plane, then the 3-plane $H$ and the line $L$ must be split. We define the subsets of ${\mathbb{X}}_{B}$ as follows ${\mathbb{X}}_{H}:={\mathbb{X}}_{B}\cap H$ and ${\mathbb{X}}_{L}:={\mathbb{X}}_{B}\cap L$. Remark that ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$. By Proposition 2.8(a), two subsets ${\mathbb{X}}_{H}$ and ${\mathbb{X}}_{L}$ also have $\mathop{\rm CBP}\nolimits(r-1)$. By Corollary 2.4, we have $|{\mathbb{X}}_{L}|\geq r+1$. Observe that ${\mathbb{X}}_{H}$ may lie on a plane configuration of dimension 2, but it does not lie on a line. By Proposition 3.1(a), the cardinality of ${\mathbb{X}}_{H}$ is bounded below as

### v2

@@ Abstract.
In this paper, we study the geometric configurations of a finite set of points having the Cayley-Bacharach property in the $n$-dimensional projective space ${\mathbb{P}}^{n}$. Our main contribution is the proof of the Levinson-Ullery conjecture for the previously unsolved case where $d=4$ and $r\geq 1$.
@@ 1. Introduction
Inspired by this result, in [16], Levinson and Ullery introduced the notion of a plane configuration to describe the geometry of $\mathbb{X}$} and proposed the following conjecture. Here a plane configuration $\mathcal{P}$ is a union of distinct positive-dimensional linear spaces $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ in ${\mathbb{P}}^{n}$, its dimension is $\dim(\mathcal{P})=\sum_{i=1}^{k}\dim(\mathcal{P}_{i})$ and its length is $l(\mathcal{P})=k$.
@@ 2. Preliminary
    The image of the vanishing ideal $I_{\mathbb{Y}}\subseteq P$ of $\mathbb{Y}$} in $R=P/I_{\mathbb{X}}$ is denoted by $I_{{\mathbb{Y}}/{\mathbb{X}}}$. Its initial degree is $\alpha_{{\mathbb{Y}}/{\mathbb{X}}}:=\min\{k\in{\mathbb{N}}\mid(I_{{\mathbb{Y}}/{\mathbb{X}}})_{k}\neq 0\}$. Then $\alpha_{{\mathbb{Y}}/{\mathbb{X}}}\leq r_{\mathbb{X}}$ and the Hilbert function of $\mathbb{Y}$} satisfies
@@
Now we recall from [16, Definition 2.1] the following notion of plane configurations.
@@
Note that a 0-dimensional plane configuration is empty, a 1-dimensional plane configuration is a single line, and a 2-dimensional plane configuration is either two lines or one 2-plane. Here a $d$-plane refers to a $d$-dimensional linear space. Also, the linear span of a plane configuration $\mathcal{P}=\bigcup_{i=1}^{k}\mathcal{P}_{i}$ is denoted by $\mathop{\rm span}\nolimits(\mathcal{P})$ or $\mathop{\rm span}\nolimits(\mathcal{P}_{i}\mid i=1,\dots,k)$.
@@ 3. Proof of main theorem
    AA and $L$ are split. In this case, ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{B}$ have $\mathop{\rm CBP}\nolimits(r)$ by Proposition 2.8(a). Due to Corollary 2.4 and (a), the lower bounds for the cardinalities of ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{B}$ are given by
@@
    This implies $(d+1)r+4\leq|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|=|{\mathbb{X}}|\leq(d+1)r+1$, which is impossible.
@@
    AA and $L$ intersect at a point, say $p$. By Lemma 2.8(c), at least one of ${\mathbb{X}}_{A}$ and ${\mathbb{X}}_{A}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r).$ If ${\mathbb{X}}_{A}$ has $\mathop{\rm CBP}\nolimits(r)$, then claim (a) yields that $|{\mathbb{X}}_{A}|\geq dr+2$. When ${\mathbb{X}}_{A}\cup\{p\}$ has $\mathop{\rm CBP}\nolimits(r),$ then $|{\mathbb{X}}_{A}\cup\{p\}|\geq dr+2$ by (a). In these both cases, we always have $|{\mathbb{X}}_{A}|\geq dr+1$. On the other hand, the set ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$, and Corollary 2.4 shows that $|{\mathbb{X}}_{B}|\geq r+1$. Hence, we obtain $|{\mathbb{X}}|=|{\mathbb{X}}_{A}|+|{\mathbb{X}}_{B}|\geq(dr+1)+(r+1)\geq(d+1)r+2$, a contradiction.
@@ 1. Case 1:
        If $|{\mathbb{X}}_{B}|=2r\leq 3(r-1)+1$, then ${\mathbb{X}}_{B}$ lies on a plane configuration of dimension 2 by the inductive hypothesis and $|{\mathbb{X}}_{A}|=3r+1$. When ${\mathbb{X}}_{B}$ lies on a line, as above, the set $\mathbb{X}$} lies on a plane configuration of dimension 4 and we are done. If ${\mathbb{X}}_{B}$ lies on a union of two split lines, then each of these lines contains at least $r+1$ points of ${\mathbb{X}}_{B}$ by Proposition 2.8(a) and Corollary 2.4, and so $|{\mathbb{X}}_{B}|\geq 2r+2$, which is impossible. Here we remark that a skew plane configuration of two lines agrees with a split plane configuration of two lines. Now we consider the last subcase that the set ${\mathbb{X}}_{B}$ lies on a $2$-plane $H^{\prime}$. If $H^{\prime}\cap H\neq\emptyset$ or $H^{\prime}\cap L\neq\emptyset$ then the pair $(H,H^{\prime})$ defines a 4-plane containing at least $4r$ points of $\mathbb{X}$} or the pair $(L,H^{\prime})$ and a point in ${\mathbb{X}}_{A}\cap H$ define a 4-plane containing at least $3r+2$ points of $\mathbb{X}$}, which contradicts to the choice of the 4-plane $A$. Hence we must have $H^{\prime}\cap H=H^{\prime}\cap L=\emptyset$. In particular, $\mathbb{X}$} is contained in a skew plane configuration $\mathcal{P}=H\cup L\cup H^{\prime}$ of length $\ell(\mathcal{P})=3<r$. Since $\mathbb{X}$} has $\mathop{\rm CBP}\nolimits(r)$, an application of Proposition 2.8(b) yields that the line $L$ contains at least $r+2$ points of $\mathbb{X}$}. It follows that
@@ 1. Case 2:
    Since ${\mathbb{X}}_{B}$ is not contained in a $4$-plane, then the 3-plane $H$ and the line $L$ must be split. We define two subsets of ${\mathbb{X}}_{B}$ by ${\mathbb{X}}_{H}:={\mathbb{X}}_{B}\cap H$ and ${\mathbb{X}}_{L}:={\mathbb{X}}_{B}\cap L$. Remark that ${\mathbb{X}}_{B}$ has $\mathop{\rm CBP}\nolimits(r-1)$. By Proposition 2.8(a), two subsets ${\mathbb{X}}_{H}$ and ${\mathbb{X}}_{L}$ also have $\mathop{\rm CBP}\nolimits(r-1)$. By Corollary 2.4, we have $|{\mathbb{X}}_{L}|\geq r+1$. Observe that ${\mathbb{X}}_{H}$ may lie on a plane configuration of dimension 2, but it does not lie on a line. By Proposition 3.1(a), the cardinality of ${\mathbb{X}}_{H}$ is bounded below as